ELECTRONIC GAS PHASE SPECTRA OF LARGER POLYACETYLENE AND PROTONATED POLY-ACETYLENE CATIONS

ANATOLY DZHONSON, \underline{EVAN B. JOCHNOWITZ}, EUNSOOK KIM, and JOHN P. MAIER, Department of Chemistry, University of Basel, Klingelbergstrasse 80, CH-4056 Basel, Switzerland.

A new approach has been developed for the purpose of measuring electronic transitions to bound excited states for cations that have been collisionally relaxed to low vibrational and rotational temperatures (∼25 K). This has been used to obtain the first gas phase spectra of three new polyacetylene cation chains (HC_{12}H^{+}, HC_{14}H^{+}, and HC_{16}H^{+}) as well as two different protonated species (HC_{6}H_{2}^{+} and HC_{8}H_{2}^{+}) using a two-color ion-photodissociation approach applied to ions held in a 22-pole trap. Specifically, the origin bands of the Ã ^{2}Π - X̃ ^{2}Π electronic transitions of HC_{12}H^{+}, HC_{14}H^{+}, and HC_{16}H^{+} were located at 10811, 9333, and 8739 cm^{−1}, respectively, while the B̃ ^{1}A_{1} - X̃ ^{1}A_{1} transitions of HC_{6}H_{2}^{+} and HC_{8}H_{2}^{+} (C_{2v} geometry) were recorded at 26404.0 and 21399.8 cm^{−1}.